OpenStudy (anonymous):

Graph g(x)=sqr x-3 -1

OpenStudy (anonymous):

\[\sqrt{x-3}-1\]

OpenStudy (anonymous):

\[y=\sqrt{x-3}-1\] \[y+1=\sqrt{x-3}\]\[y^2+2y+1=x-3\] \[x=y^2+2y+4\] Which is a quadratic equation, thus your graph will be a curve plug in arbitary values of x or y and calculate the other variable and plot the graph

OpenStudy (anonymous):

example of plugging in please

OpenStudy (anonymous):

example... x=4, y=-2

OpenStudy (anonymous):

hmm sorry can you write it out

OpenStudy (anonymous):

\[x=y^2+2y+4\] \[y=-2\]\[x=(-2)^2+2(-2)+4=4-4+4=4\]

OpenStudy (anonymous):

ok thanks